OpenStudy (anonymous):

The product of 3 and a number is 12 more than the sum of the number and 26. What is the number?

OpenStudy (anonymous):

@DanJS help me pleaseeee

OpenStudy (anonymous):

ineed help

OpenStudy (anonymous):

there is 226 members online and nobody can help

OpenStudy (anonymous):

medal if helped

OpenStudy (anonymous):

heeeeeeeeeeeeeeeeeeloooooooooooooo

OpenStudy (anonymous):

pls

OpenStudy (anonymous):

some one this is the last question pls

OpenStudy (danjs):

the product of 3 and a number (3*x) is = 12 more than the sum of the number and 26 (x+26) + 12

OpenStudy (danjs):

3x = x+26+12

OpenStudy (danjs):

3x = x + 38 2x = 38 x = 19